OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle. sin 8x cos x - cos 8x sin x

OpenStudy (anonymous):

sin(7x)

OpenStudy (rock_mit182):

hi, remember that: \[\sin (a-b)= \sin(a)\cos(b)-\sin(b)\cos(a) \]

OpenStudy (rock_mit182):

\[\sin(8x-x)= \sin (7x)\]